OpenStudy (kainui):

f(y+x)+f(y-x)=2f(x)f(y) what's f(x)?

OpenStudy (anonymous):

Represents a function in respect to a variable, in that case x

OpenStudy (kainui):

To make my question more clear, what is the function that satisfies this functional equation...

OpenStudy (alekos):

I think he's trying to be funny?

OpenStudy (kainui):

lol beats me.

OpenStudy (anonymous):

Obviously distribute the f first...noobs.

OpenStudy (kainui):

f(y+x)+f(y-x)=2f(x)f(y) fy+fx+fy-fx=2f^2xy 2fy=2f^2xy 1=fx f=1/x solved! ty @iambatman

OpenStudy (anonymous):

loooool, np :)